OpenStudy (anonymous):

Solve. 5y - 10 = -25.

OpenStudy (anonymous):

divide both sides by 5 then it reduces to y-2=-5 hence y=2-5=-3.@Lion_King123

OpenStudy (anonymous):

Thank you so Much

OpenStudy (anonymous):

A)y = 3 B)y = 7 C) y = -3 D) y = -7

OpenStudy (anonymous):

C obviously.:)

OpenStudy (anonymous):

lol Thanks again XD